A new machine has just been installed to cut and rough-shape large slugs. The slugs are then transferred to a precision grinder. One of the critical measurements is the outside diameter. The quality control inspector randomly selected five slugs each half-hour, measured the outside diameter, and recorded the results. The measurements (in millimeters) for the period 8:00 a.m. to 10:30 a.m. follow.

Outside Diameter (millimeters)

Time

1

2

3

4

5

8:00

87.3

87.5

87.9

87.2

87.0

8:30

86.9

88.5

87.6

87.5

87.4

9:00

87.7

88.5

86.7

87.5

88.2

9:30

86.0

88.0

87.2

87.6

87.1

10:00

86.9

87.3

87.2

87.3

87.2

10:30

88.2

86.5

87.3

87.2

87.8

  1. Determine the control limits for the mean and the range.(Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 2 decimal places.)

Mean of the sample means

=

LCL

=

UCL

=

Mean of the sample ranges

=

LCL

=

UCL

=

  1. c-1. Are any means out of control?

Yes

No

  1. c-2. Are any ranges out of control?

Yes

No
Answer #1

(a)

Mean of the sample means:  LCL =86.60, UCL=88.22

Mean of the sample ranges : LCL =0, UCL=2.96

here \bar{x}=87.41 and \bar{R}=1.40

x-bar chart (for n=5, A2=0.577)

LCL=\bar{x}-A2*\bar{R}=87.41-0.577*1.40=86.60

UCL=\bar{x}+A2*\bar{R}=87.41+0.577*1.40=88.22

R-bar chart (for n=5, D3=0 , D4=2.114)

LCL=D3*\bar{R}=0.*1.40=0

UCL=D4*\bar{R}=2.114*1.40=2.96

Time/sample 1 2 3 4 5 mean range
8:00 87.3 87.5 87.9 87.2 87 87.38 0.90
8:30 86.9 88.5 87.6 87.5 87.4 87.58 1.60
9:00 87.7 88.5 86.7 87.5 88.2 87.72 1.80
9:30 86 88 87.2 87.6 87.1 87.18 2.00
10:00 86.9 87.3 87.2 87.3 87.2 87.18 0.40
10:30 88.2 86.5 87.3 87.2 87.8 87.40 1.70
mean= 87.41 1.40

(c1) NO, any means is not out of control

since all the means are within the LCL and UCL, so system is not out of contorll

x-bar chart mean mean CL LCL *UCL 1 2 3 4 5 6 sample

(c2)NO, any range is not out of control

since all the means are within the LCL and UCL, so system is not out of contorll

R-control chart Range range CL LCL UCL 1 2 3 4 5 6 7 sample